The average mark of candidates in an aptitude test was 128.5 with a standard deviation of 8.2. Three scores extracted from the test are; 148, 102, 152. What is the average of the extracted scores that are extreme values (outliers)?
1
Expert's answer
2020-06-02T18:29:29-0400

"\text{According to the Empirical rule any}\\
\text{ value outside two standard deviations}\\
\text{ِ of the mean is considered an outlier}\\
upper\; limit=128.5+2(8.2)=144.9,\\
lower \;limit=128.5-2(8.2)=112.1\\
\text{the range of two standard deviation}\\
\text{ of the mean is (112.1,144.9)}\\
\text{So all the values are outliers}\\
\text{the average of the extracted scores that }\\
\text{are extreme values (outliers)}
\\=\frac{148+102+152}{3}=134"